HETAV PATEL

Grade 11,

A servo voltage stabilizer restricts the voltage output to 220 V ± 1%. If an electric bulb rated at 220 V, 100 W is connected to it, what will be the minimum and maximum power consumed by it?

A servo voltage stabilizer restricts the voltage output to 220 V ± 1%. If an electric bulb rated at 220 V, 100 W is connected to it, what will be the minimum and maximum power consumed by it?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. Output voltage = 220 ± 1% 1% of 220 V = 2.2 v The resistance of bulb R = V^2/P = (220)^2/100 = 484 Ω (a) For minimum power consumed V base 1 = 2220 – 1% = 220 – 2.2 = 217.8 ∴ i = V base 1/R = 217.8/484 = 0.45 A Power consumed = i * V base 1 = 0.45 * 217.8 = 98.01 W (b) for maximum power consumed V base 2 = 220 + 1% = 220 + 2.2 = 222.2 ∴ i = V base 2/R = 222.2/484 = 0.459 Power consumed = i * V base 2 = 0.459 * 222.2 = 102 W